# MSI-NeRF: Linking Omni-Depth with View Synthesis through Multi-Sphere Image aided Generalizable Neural Radiance Field

### v1

@@ III-C Multi-Task Semi-Self-Supervision
If only input fisheye images are used for supervision, the network will suffer from serious over-fitting and lose the ability to predict scene information from other perspectives. However, after adding depth supervision, correct geometry guidance can be added, and such over-fitting can be well-constrained. In this way, the correct scene appearance can be learned. The structure of explicit MSI volumes and implicit NeRF rendering also works as a perfect geometry proxy, keeping the network away from over-fitting.
IV Experiments
IV-A Implementation Details
We implement our method using the Pytorch framework. Training and validation are all carried out on a GPU server with four RTX 2080Ti. During optimization, we use ADAM [24] optimizer with a learning rate of $3\times 10^{-4}$. The loss weight of depth is set to $\lambda_{d}=5$. We first train the base model on a generic dataset for $30$ epochs and then fine-tune it on another specific dataset for $10$. We set the inverse depth layer number to $64$ and the range of layer value to $\left[0,2\right]m^{-1}$. Due to limited GPU memory, after inference of geometry and appearance volumes, we randomly sample $8192$ rays on fisheyes and $8192$ rays on the omnidirectional depth map, expressed as $|\mathcal{R}_{p}|=|\mathcal{R}_{f}|=8192$. Then, volume rendering is done on these rays for further color and depth supervision.
For comparison, we keep the network capacity and hyperparameters of baseline methods consistent with our implementation. The same training strategy is used to match the data format. In particular, we add a depth regressor to the output module of MatryODShka [18] so that it can be trained using depth data.
IV-B Datasets
@@
TABLE I: Quantitative evaluation of omnidirectional depth estimation.
@@
IV-C Depth Evaluation
We evaluate the omnidirectional depth estimation performance on a validation set extracted from the OmniHouse and OmniThings. Original and fine-tuned models are used for different datasets, respectively. The qualitative and quantitative evaluation results are shown in Fig. 4 and Tab. I. We use inverse depth value as our comparison metric, and the unit is set to $m^{-1}$. The qualifier ”$>k$” represents the pixel ratio whose error is larger than threshold $k$.
@@ IV-C Depth Evaluation
IV-D Novel View Synthesis Evaluation
To test the novel view synthesis ability, we conduct comparative experiments against MatryODShka [18] and NeRF-360-NGP [6, 16, 20]. For MatryODShka, we replace the original omni-stereo with fisheye images to align the input format. For NeRF-360-NGP, we expand the NeRF method to $360^{\circ}$ using spatial parameterization from Mip-NeRF 360 [16] and achieve efficient training through Instant-NGP [20].
We show our qualitative and quantitative results in Fig. 5 and Tab. II. It can be seen that the color blending design in MatryODShka generates ghosting effect when trained with only depth ground truth. NeRF-360-NGP fails to generate reasonable novel views with limited input images and shared viewing angles. Our method exceeds other generalizable and scene-specific methods in terms of image quality. It achieves high accuracy on novel view synthesis and shows excellent generalization ability on cross-dataset validation.
TABLE II: Quantitative evaluation of panoramic novel view synthesis.
@@ IV-D Novel View Synthesis Evaluation
IV-E Ablation Studies
Our generalizable neural rendering method does not follow traditional novel view synthesis process. It is mainly reflected in three key designs: additional color projection, radiance field rendering, and separated geometry and appearance decoder. We conduct ablation studies to demonstrate the effect of these designs by removing them individually. The results are shown in Fig. 7 and Tab. III.
@@ IV-E Ablation Studies
IV-F Real-World Experiment
@@
Figure 6: (a): Our self-made multi-fisheye camera rig system for real-world data capture. It comprises $4$ USB fisheye cameras with $220^{\circ}$ FoV arranged on a circle with a diameter of $0.4m$. (b): The captured source views, the generated omnidirectional depth map, and the synthesized image in the real-world experiment.
@@
To test our approach’s generalizability and real-world performance, we constructed a self-collected dataset using a multi-fisheye camera rig with similar settings to the synthetic datasets. The data capture system, fisheye image samples, and the generated results can be seen in Fig. 6. From the generated depth and novel view images, we can see that our method can be well generalized from sim to real. The real-world experiment also demonstrates the practical application value of this method for virtual reality video shooting and 3D scene reconstruction in real-world scenes.
V Conclusion
@@
This paper explores the problem of generating a panoramic scene representation using only four fisheye images as input. To solve this, our presented MSI-NeRF combines the advantage of multi-sphere image representation with volume rendering to construct a generalizable NeRF. From such representation, accurate omnidirectional depth estimation and high-quality novel view synthesis can be conducted. Our proposed semi-self-supervision strategy allows the network to be trained using only depth ground truth. Furthermore, generalizability also enables it to reconstruct unknown scenes efficiently, which is of great significance for robotics and virtual reality applications. As a result, it outperforms baseline methods in both image quality and feasibility.
Figure 7: Qualitative results of our ablation study. We generate novel view images and depth maps of our method without key designs and compare them with our full model. Ours full model generates the finest novel view images while maintaining the depth map quality.
TABLE III: Quantitative evaluation of ablation study. We compare both omnidirectional depth and novel view synthesis metrics.
Method $>$ 0.1 $>$ 0.3 $>$ 0.5 MAE RMSE PSNR SSIM LPIPS
w/o Color Input 1.021 0.140 0.030 0.012 0.024 32.080 0.783 0.287
w/o Radiance Field 1.152 0.308 0.145 0.012 0.023 29.620 0.553 0.432
w/o Separated Decoder 1.396 0.266 0.079 0.013 0.026 37.200 0.953 0.088
Ours Full Model 1.255 0.180 0.033 0.011 0.024 37.302 0.957 0.077

### v2

@@ 3.3 Multi-Task Supervision
The depth supervision can provide a useful geometry guidance to the multi-sphere images, which also shares knowledge to appearance network and benefit color image generation. If only source fisheye images are used for training, turning the method into fully self-supervised, the network will suffer from insufficient overlapping and bad scene structure prediction. Adding depth supervision greatly constrained such problems.
4 Experiments
@@
Table 1: Quantitative evaluation result of depth estimation.
4.1 Implementation Details
We implement our method using the Pytorch framework. Training and validation are all carried out on a GPU server with four RTX 2080Ti. During optimization, we use ADAM [21] optimizer with a learning rate of $3\times 10^{-4}$. The loss weight of depth is set to $\lambda_{d}=5$. We first train the base model on OmniThings dataset for $30$ epochs and then fine-tune it on OmniHouse dataset for $10$ [38]. We set the inverse depth layer number to $64$ and the range of layer value to $\left[0,2\right]m^{-1}$. Due to limited GPU memory, after inference of geometry and appearance volumes, we randomly sample $8192$ rays on fisheyes and $8192$ rays on the omnidirectional depth map, expressed as $|\mathcal{R}_{p}|=|\mathcal{R}_{f}|=8192$. Then, volume rendering is done on these rays for further color and depth supervision.
For comparison, baseline methods are retrained on the same datasets using the same training strategy. We also keep the network capacity and hyperparameters of them consistent with our implementation.
4.2 Datasets
@@
4.3 Depth Evaluation
We evaluate the omnidirectional depth estimation performance on a validation set extracted from the OmniHouse and OmniThings. Original and fine-tuned models are used for different datasets, respectively. The qualitative and quantitative evaluation results are shown in Fig. 3 and Tab. 1. We use inverse depth value as our comparison metric, and the unit is set to $m^{-1}$. The qualifier ”$>k$” represents the pixel ratio whose MAE is larger than threshold $k$.
@@ 4.3 Depth Evaluation
4.4 Novel View Synthesis Evaluation
To test the novel view synthesis ability, we conduct comparative experiments against MatryODShka [1] and NeRF-360-NGP [29, 3, 31]. Since MatryODShka uses binocular omnidirectional stereo (ODS) images as input which is also originated from fisheye cameras, we re-implement it using our fisheye input setting to align the input format. In particular, we use sphere sweep method directly on the input fisheye images to generate four sphere sweep volumes (SSV) and pass them to the network to get the blending weights. We also add a depth renderer so that it can be trained using depth ground-truth only. As for NeRF-360-NGP, we expand the NeRF [29] method to $360^{\circ}$ using spatial parameterization from Mip-NeRF 360 [3] and achieve efficient training through Instant-NGP [31].
We show our qualitative and quantitative results in Fig. 5 and Tab. 2. It can be seen that the color blending design in MatryODShka generates ghosting effect when trained with only depth ground truth. NeRF-360-NGP fails to generate reasonable novel views with limited input images and shared viewing angles. Our method exceeds other generalizable and scene-specific methods in terms of image quality. It achieves high accuracy on novel view synthesis and shows excellent generalization ability on cross-dataset validation.
@@ 4.4 Novel View Synthesis Evaluation
Table 2: Quantitative evaluation result of novel view synthesis.
Figure 6: Qualitative results of our ablation studies. We generate novel view images of our method without key designs and compare them with our full model. Ours full model generates the novel view images with the best overall quality and finest details.
4.5 Ablation Studies
Our generalizable neural rendering method does not follow traditional novel view synthesis process. It is mainly reflected in three key designs: additional color projection, radiance field rendering, and separated geometry and appearance decoder. We conduct ablation studies to demonstrate the effect of these designs by removing them individually. The results are shown in Fig. 6 and Tab. 3.
@@ 4.5 Ablation Studies
Figure 7: Visualization of our method with only color supervision
We also test our method with only color image for supervision (self-supervised), and the results are shown in Fig. 7. Although the structure of explicit MSI volumes along with implicit NeRF rendering works as geometry proxy and reasonable rendering results can be generated, they are of severe artifacts which are far from being able to be applied in upper-level applications.
Method PSNR SSIM LPIPS
w/o Color Input 32.080 0.783 0.287
w/o Radiance Field 29.620 0.553 0.432
w/o Separated Decoder 37.200 0.953 0.088
Ours Full Model 37.302 0.957 0.077
Table 3: Quantitative evaluation result of ablation studies.
4.6 Real-world Experiment
@@
Figure 8: (a): Our self-made multi-fisheye camera rig system for real-world data capture. It comprises $4$ fisheye cameras with $220^{\circ}$ FoV arranged on a circle with a diameter of $0.4m$. (b): The captured source views, the generated omnidirectional depth map, and the synthesized image in the real-world experiment.
@@
To test our approach’s generalizability and real-world performance, we constructed a self-collected dataset using a multi-fisheye camera rig with similar settings to the synthetic datasets. The data capture system, fisheye image samples, and the generated results can be seen in Fig. 8. From the generated depth and novel view images, we can see that our method can be well generalized from sim to real. The real-world experiment also demonstrates the practical application value of this method for virtual reality video shooting and 3D scene reconstruction in real-world scenes. However, our method has limitations because it requires accurate camera intrinsic and extrinsic parameters and calibration errors may cause blurring and artifacts in color and depth images in real-world scenes.
5 Conclusion
@@
This paper explores the problem of generating a panoramic scene representation using only four fisheye images as input. To solve this, our presented MSI-NeRF combines the advantage of multi-sphere image representation with volume rendering to construct a generalizable NeRF. From such representation, accurate omnidirectional depth estimation and high-quality novel view synthesis can be conducted efficiently, with only a single network inference. As a result, our method outperforms baseline methods in both image quality and feasibility, which is of great significance for virtual reality and robotics applications.
6 Additional Implementation Details
Since MatryODShka [1] has different input and output format to our method, to achieve fair comparison, we re-implement MatryODShka and match its data format with ours. MatryODShka uses two omnidirectional stereo (ODS) images as input and sphere sweep the images as a sphere sweep volume (SSV). We find that four fisheye images can provide larger FoV than the original two ODS images, so we directly use fisheye images to construct the SSV. For the network MatryODShka used to process SSV, we use the same 3D CNN structure as our MSI decoder by just changing the input and output channel number. We modify the output of the network to generate a blending weight using soft max layer. Alpha blending is then conducted and the final color and depth images can be rendered.
7 Network Architecture
We basically follow the CNN structure proposed in [38] and modify the output format to connect it with our NeRF [29] MLP. The CNN network channel is declined to save GPU memory for the additional MLP and neural rendering. The input of the network is four fisheye images. We use the residual convolution blocks for the image feature extraction, and the dilated convolution for the larger receptive field. The output feature map size is half of the input image. We use a shallow MLP that takes coordinate, direction, projected color and interpolated feature vector as input. The output color and occupancy are then rendered into color and depth map.
8 Derivation Details
We give a detailed derivation of Equation (6) in the main body. Given the ray origin $\mathbf{r}_{o}=(o_{x},o_{y},o_{z})^{T}$ and ray direction $\mathbf{r}_{d}=(d_{x},d_{y},d_{z})^{T}$, the equation of the ray can be given as:
$\mathbf{r}(z)=\mathbf{r}_{o}+z\mathbf{r}_{d},$ (12)
where z is the parameter indicating the distance along the ray, the depth. Since the sphere equation can be writen as $x^{2}+y^{2}+z^{2}=d^{2}$, we substitute $\mathbf{r}(z)$ into the sphere equation:
$\left(r_{o_{x}}+zr_{d_{x}}\right)^{2}+\left(r_{o_{y}}+zr_{d_{y}}\right)^{2}+\left(r_{o_{z}}+zr_{d_{z}}\right)^{2}=d^{2}.$ (13)
We can simplify it to:
$r_{o_{x}}^{2}+r_{o_{y}}^{2}+r_{o_{z}}^{2}+2z\left(r_{o_{x}}r_{d_{x}}+r_{o_{y}}r_{d_{y}}+r_{o_{z}}r_{d_{z}}\right)+z^{2}=d^{2}.$ (14)
From this, we can tell that the problem is to solve a quadratic equation and the coefficients of the quadratic term, linear term, constant term can be represented as:
$a=\mathbf{r}_{d}^{T}\mathbf{r}_{d},\quad b=2\mathbf{r}_{d}^{T}\mathbf{r}_{o},\quad c_{n}=\mathbf{r}_{o}^{T}\mathbf{r}_{o}-1/\left(d^{-1}\right)^{2}.$ (15)
Then the original formula can be established.
9 Application Prospects
Being able to synthesize novel views in 6DoF and estimate depth map panoramically mean a lot to VR and robotics.
Traditional VR contents are usually shot by image stitching, which is simply a stream of panorama images. This results in a limited 3DoF camera movement which can cause omitted parallax information, low immersion, and VR sickness. Instead, our method retains the parallax information within the multi-camera system and achieves 6DoF rendering. This can bring a whole new experience to VR content without changing the shooting equipment. Further more, different from the existing panoramic cameras that only support three-degree-of-freedom re-movement and editing, the video captured by this method can be moved and edited in the translation direction. This is of great significance for filming and video making.
On the robotics side, being able to produce omnidirectional depth estimation also helps a lot. It bring perception in autonomous robots to a new level. Compared to LiDAR, it is much cheaper, with wider FoV, and denser points per frame. The novel view synthesis ability also makes driving cars or drones under a third person’s view possible, which benefits a lot in teleoperation and navigation.
10 Limitations
Limited by the memory size of the GPU device, our network capacity is restricted and the resolution of the generated color and depth map is set to $512\times 256$. Also, due to the limited computation performance, our method can only run at around three frames per second. These limitations makes our method hard to applied in applications with strong real-time requirements.
At the same time, our method also requires accurate extrinsic and intrinsic calibration. These parameters are used in the warping process that turns fisheye projection into MSI representation, which is essential to the final result. In our real-world experiment, the generated color and depth image both suffer from poorly calibrated cameras, which results in blurry and artifacts. The quality of real-world generation will degrade compared to the results from synthetic data.
Due to the usage of MSI, the free roaming of the camera is limited inside the smallest sphere. This way, the setting of the smallest radius can be a trade of between the nearest collision limit and the virtual camera moving range. A bigger inner sphere can lead to a wider moving range, while may cause objects close to camera being cut off. This trade off needs careful tuning when faced with different scenarios.
11 Additional Experiments Result
Here, we show the additional generated depth estimation and novel view synthesis results on the OmniHouse dataset. They are shown in Fig. 9. More visualization results can be seen in the video in the supplementary files.
Figure 9: Additional experiment results.